The mean distance of two planets A and B from sun is 2 and 4 times distance of Earth from the sun respectively . Find the ratio of time taken by two planets to make one revolution around the sun ?

Grade:9

1 Answers

Arun
6 years ago
As we know,
From kepler law,
T^2= k* r^3
 
Hence
(T1/T2)^2 = (r1/r2)^3
(T1/T2)^2  = (2r/4r)^3
T1/T2 = \sqrt(1/8)
T1/T2 = 1/2\sqrt2